If A

 and B

 are vector fields, prove the following:


∇(A⋅B)=(B⋅∇)A+(A⋅∇)B+B×(∇×A)+A×(∇×B).

Submit

Some problems have options such as save, reset, hints, or show answer. These options follow the Submit button.



1
Expert's answer
2021-12-06T16:08:04-0500

Consider,

"\nabla(\vec{A}\cdot\vec{B})=\sum\vec{i}\frac{d}{dx}(\vec{A}\cdot\vec{B})=\sum\vec{i}(\frac{d\vec{A}}{dx}\cdot\vec{B}+\vec{A}\cdot\frac{d\vec{B}}{dx})"

"=\sum\vec{i}(\frac{d\vec{A}}{dx}\cdot\vec{B})+\sum(\vec{A}\cdot\frac{d\vec{B}}{dx})" ..............(i)

Now "\vec{B}(\vec{i}*\frac{d\vec{A}}{dx})=(\vec{B}\cdot\frac{d\vec{A}}{dx})\vec{i}-(\vec{B}\cdot\vec{i})\frac{d\vec{A}}{dx}"

"=(\vec{B}\cdot\frac{d\vec{A}}{dx})\vec{i}=\vec{B}*(\vec{i}*\frac{d\vec{A}}{dx})+(\vec{B}\cdot\vec{i})\frac{d\vec{A}}{dx}"

"\therefore \sum(\vec{B}\cdot\frac{d\vec{A}}{dx})\vec{i}=\vec{B}*(\vec{i}*\frac{d\vec{A}}{dx})+(\vec{B}\cdot\vec{i})\frac{d\vec{A}}{dx}"

"=\vec{B}*\sum(\vec{i}*\frac{d\vec{A}}{dx})+(\vec{B}\cdot\sum\vec{i}\frac{d}{dx})\vec{A}"

"=\vec{B}*(\nabla*\vec{A})+(\vec{B}\cdot\nabla)\vec{A}"

"\therefore\sum\vec{i}(\frac{d\vec{A}}{dx}\cdot\vec{B})=\vec{B}*(\nabla*\vec{A})+(\vec{B}\cdot\nabla)\vec{A}............(ii)"

Similarly, if we interchange the role of "\vec{A}" and "\vec{B}" we can prove;

"\sum\vec{i}(\vec{A}\cdot\frac{d\vec{B}}{dx})=\vec{A}*(\nabla*\vec{B})+(\vec{A}\cdot\nabla)\vec{B}............(iii)"

Substituting (ii) and (iii) in (i), we get:

"\nabla(\vec{A}\cdot\vec{B})=(\vec{B}\cdot\nabla)\vec{A}+(\vec{A}\cdot\nabla)\vec{B}+\vec{B}*(\nabla*\vec{A})+\vec{A}*(\nabla*\vec{A})"